OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 1) greater than or equal to -12? x greater than or equal to -5 x less than or greater to -5 -5 less than or greater to x A number line with a closed circle on negative 5 and shading to the right @jziggy

OpenStudy (jziggy):

2(x-1) >= -12 x - 1 >= -6 x >= -5

OpenStudy (jziggy):

So.... 1. is obviously a way to represent it. Since 2 is the opposite of that, it's obviously not true. 3. doesn't seem to mean anything..... not sure how a number can be less than and greater than a variable.... do you mean -5 is less than *or equal to* x Finally 4 is also a way to represent it as long as the circle is shaded to include the equal to part.

OpenStudy (anonymous):

Thank you! Can you help with more?

OpenStudy (jziggy):

Sure

OpenStudy (anonymous):

The table shows the solution to the equation |2x - 5| - 2 = 3: Step 1 |2x - 5| = 3 + 2 Step 2 |2x - 5| = 5 Step 3 2x - 5 = 5 or 2x - 5 = -5 Step 4 2x = 10 or 2x = 0 Step 5 x = 5 or x = 0 Which is the first incorrect step? Step 3 Step 4 Step 5 Solution is correct

OpenStudy (jziggy):

Well, first check if the solution is correct. try plugging in the answers in step 5 to the original equation.

OpenStudy (anonymous):

I'm really not good at solving this kind of stuff. I'm just really trying to finish these tests fast. :/

OpenStudy (jziggy):

|2x-5| = 5 Solution 1: |2(5) - 5| = 5 Solution 2: |2(0) - 5| = 5 Now just simplify to see if it works.

OpenStudy (anonymous):

How would I simplify that? :/

OpenStudy (jziggy):

Well, start by multiplying the 2 with the number in the parentheses.

OpenStudy (anonymous):

2*(5) = 10 - 5 | = 5?

OpenStudy (jziggy):

yeah, so solution 1 is correct now let's check solution 2.

OpenStudy (jziggy):

|2(0)-5| = 5

OpenStudy (anonymous):

Solution 2 isn't correct, so the answer is B?

OpenStudy (anonymous):

No wait sorry

OpenStudy (anonymous):

Was looking at a different problem aha,

OpenStudy (jziggy):

| is an absolute value bar so when you remove them the output is always positive.

OpenStudy (jziggy):

|0-5| = 5 |-5| = 5 5 = 5

OpenStudy (jziggy):

Since both of those were true, the solution is correct.

OpenStudy (anonymous):

Oh! Okay thank you.

OpenStudy (jziggy):

Sorry, I have to get some food now, good luck with the rest :)